OpenStudy (anonymous):

A swimming pool is 60 ft. long and 35 ft. wide. A model of the pool is 9 in. long. How wide is the model?

OpenStudy (anonymous):

5.25 in

OpenStudy (anonymous):

ok plz tell me how u got that?

OpenStudy (anonymous):

please go see my question http://openstudy.com/#/updates/4ea97615e4b02eee39d5a1ac

OpenStudy (anonymous):

60 ft = 9 in 1 ft = 9/60 in 35 ft = (9*35)/60 in = 5.25 in